# GoDe: Gaussians on Demand for Progressive Level of Detail and Scalable Compression

### v1

@@ Abstract
3D Gaussian Splatting enhances real-time performance in novel view synthesis by representing scenes with mixtures of Gaussians and utilizing differentiable rasterization. However, it typically requires large storage capacity and high VRAM, demanding the design of effective pruning and compression techniques. Existing methods, while effective in some scenarios, struggle with scalability and fail to adapt models based on critical factors such as computing capabilities or bandwidth, requiring to re-train the model under different configurations. In this work, we propose a novel, model-agnostic technique that organizes Gaussians into several hierarchical layers, enabling progressive Level of Detail (LoD) strategy. This method, combined with recent approach of compression of 3DGS, allows a single model to instantly scale across several compression ratios, with minimal to none impact to quality compared to a single non-scalable model and without requiring re-training. We validate our approach on typical datasets and benchmarks, showcasing low distortion and substantial gains in terms of scalability and adaptability. The source code, provided in the supplementary material, will be open-sourced upon acceptance of the paper.
@@ 1 Introduction
Figure 1: Our method constructs a progressive hierarchy of Gaussians, enabling progressive LoD. Combined with compression techniques, this allows the model to operate across several (eight in the figure) compression rates, ranging from extremely high to low. The model can dynamically adjust the rate based on criteria such as a target minimum FPS.
@@
To address this scalability issue, we propose GoDe (Gaussians on Demand), a novel approach that enables scalable compression. As shown in Fig. 1, GoDe enables a 3DGS model to scale across different compression rates, without the need for retraining specialized models. Specifically, GoDe employs a progressive Level of Detail (LoD) approach: Gaussians are organized into discrete and progressive levels, with each level adding more detail (enhancement) to the previous one. This progressive structure allows 3DGS models to be highly scalable across different scenarios. In fact, by adding or removing enhancement layers, we can easily adjust the trade-off between quality and performance. When combined with compression technique schemes, this approach enables scalable compression. Each level of detail can be seen as a distinct compression level, where higher LoDs corresponds to lower compression rate and lower LoDs corresponds to a higher compression rate. This allows for navigating the rate/distortion curve by simply adding or removing enhancement layers. To build the hierarchy, inspired by iterative gradient-informed pruning approaches, we developed an iterative gradient-informed masking methods that progressively constructs the hierarchy in a top-down manner (from high to low levels). This methods is general and compatible with popular 3DGS baselines. A fine-tuning phase ensures coherence and smoothness across the levels, while a final compression stage further reduce model size.
In summary, our main contributions are:
@@
  • A general method for progressive LoD (Sec. 3.1): Leveraging a novel gradient-based iterative masking technique, we show the applicability of this method to a variety of popular 3DGS model (Sec. 4.2) and how this approach effectively organizes primitives into a hierarchy of progressive levels with respect to other approaches (Sec. 4.4.2).
  • Smooth transition between levels (Sec. 3.2): Our fine-tuning procedure offers a simple and effective way to reduce artifacts between levels, ensuring smooth transitions in visual fidelity across multiple LoDs (Sec. 4.3) and boosting LoD quality (Sect. 4.4.3).
  • Scalable Compression (Sec. 3.3): Our method integrates and extends state-of-the-art compression techniques, enabling scalable compression and adaptive rendering, with comparable and sometimes even superior performance of non-scalable methods (Sec. 4.4.4).
@@
We extensively evaluate our approach on standard benchmark and compared it against the most recent state-of-the-art methods, showing comparable and even superior performances than non-scalable method. Our method allows models to scale across a wide range of compression levels, from low to very high, achieving up to 99.76% reduction in model size and 98.36% in the number of primitives in the highest compression rate with respect to non-compressed models, while preserving comparable full-size model quality at the lowest compression rate.
@@ 2 Background and Related Work
LoD is a well-established technique in computer graphics, used to optimize performance and improve visual quality. It comprises three main components: constructing a scene representation at various levels of detail, implementing a policy for selecting the appropriate level based on the view, and enabling seamless transitions between levels to avoid artifacts or noticeable stuttering. Recent works have integrated LoD techniques with 3DGS, mainly to reduce aliasing artifacts [38] and improve performance, especially for large-scale datasets [23]. These methods typically involve learning multiple levels of detail within the scene representation, followed by selective rendering according to heuristics like pixel coverage (i.e., the number of pixels occupied by a Gaussian projection on the image plane) and distance. Other work propose a hierarchical representation, often structured as a tree, as proposed by [24, 18]. This method dynamically implements LoD by selectively rendering Gaussians up to a specific level within the structure based on certain heuristics. In these approaches, LoD is typically defined at render time (dynamic LoD), which complicates its integration into compression contexts like ours, where a predefined, discrete set of progressive enhancement layers is essential. Additionally, these dynamic LoD methods often introduce significant overhead due to the need for tree traversal algorithms to construct levels of detail, as opposed to the simpler and faster operation of layer concatenation, which is more suitable for adaptive rendering in bandwidth-limited scenarios.
@@
Since the introduction of hybrid radiance fields models that incorporate auxiliary data structures to accelerate training and inference times, such as voxel grids [33, 5, 13, 28], research has focused on compression [8] and complexity reduction [9]. For instance, [8] demonstrates how an iterative-magnitude pruning approach is highly effective in reducing the number of parameters even in the context of radiance fields. These techniques have recently been adapted in the field of 3DGS. In fact, despite its advantages, 3DGS is limited by high storage and memory demands, reducing its application, for example, on low-end hardware. Current compression methods aim at achieving the highest quality at minimal storage size. Typically, compact representations involve post-hoc Gaussian pruning followed by compression stages, including quantization and entropy coding [1, 10, 30, 36]. Some methods operate directly during training, aiming to limit the number of Gaussians via learned masks and optimized encoding schemes [14, 6, 22, 31, 34]. Other studies enhance the representation by employing neural networks for on-the-fly Gaussian generation [24], tri-planes [35], or by using auxiliary structures such as hash grids to store spherical harmonics [6, 21], further reducing storage demands. In particular, Scaffold-GS [24] reduces storage by up to 5 times by replacing Gaussians with anchor points. These anchor points, which are significantly fewer in number, are associated with a set of features, offsets, and scalings. A set of shallow neural networks process each anchor point and generates the parameters of view-adaptive Gaussians. However, it is important to note that none of these methods are scalable in practice. While different compression rates can be achieved, they require retraining the model from scratch or performing additional fine-tuning steps for each compression level. This limitation makes it difficult to adapt these model across a broad range of context, such as low bandwidth or varying hardware capabilities. In contrast, our method incorporates a scalable compression scheme: with a single training we can instantly obtain different compression levels, effectively navigating the rate-distortion curve. Furthermore, our method matches or even surpass dedicated, non-scalable compression techniques.
@@ 3 Gaussians on demand
Figure 2: Graphical overview of our method GoDe. 1) Starting from a pre-trained model, we build a progressive hierarchy of Gaussians through iterative gradient-informed masking. 2) A fine-tuning stage further enhances the model’s quality and reduces its memory footprint by culling unnecessary SH coefficients. 3) The model is further compressed through codebook quantization.
This section presents our proposed efficient and adaptive 3D rendering method using hierarchical Gaussian representations. The proposed approach extends the standard 3D Gaussian Splatting (3DGS) framework by introducing a hierarchical representation of the scene at multiple levels of detail. The objective is to develop a single, scalable model that can dynamically adapt to various criteria, such as computational resources or available bandwidth. We present a novel algorithm that implicitly constructs this hierarchy, enabling a single 3DGS model to learn multiple levels of detail. We first describe how to build a progressive hierarchy of Gaussians from a pre-trained 3DGS model (Sec. 3.1) . Then, we show how to train and fine-tune Gaussians in different levels of detail (Sec. 3.2). Finally, we describe our scalable compression and adaptive rendering steps (Sec. 3.3).
@@ 3.1 Hierarchy of Gaussians
$\mathcal{G}=(\mathcal{G}_{0},~{}\mathcal{E}_{1},\dots,\mathcal{E}_{l},\dots,\mathcal{E}_{L-1}),$ (1)
@@
$\mathcal{G}_{0}\cap\mathcal{E}_{1},\dots,\cap~{}\mathcal{E}_{l},\dots,\cap~{}\mathcal{E}_{L-1}=\emptyset.$ (3)
@@
and $c_{\text{min}}=|\mathcal{G}_{0}|$ is the number of Gaussians at the first LoD level (chosen as desired). We choose this progression since the rate-distortion functions generally follow a logarithmic law (see the theoretical rate-distortion curve for a Gaussian source [7]).
To organize the Gaussians of the model into a hierarchical structure, we employ an iterative gradient-informed masking process. This approach follows a top-down strategy, starting with the construction of the higher enhancement levels first, namely, progressively building the model’s hierarchy from layer $l=L-1$ to layer $l=1$. The gradient of each parameter in the Gaussians serves as a critical and general indicator of its relevance to the overall model. A Gaussian with a low gradient indicates minimal contribution to the loss function. Conversely, a high gradient suggests that the Gaussian is essential and has a significant impact on the loss. We can then leverage gradient information to build a hierarchy. Specifically, to create the enhancement set $\mathcal{E}_{l}$, we accumulate gradients on $\mathcal{G}_{l}$ using the training set in order to select the $k_{l}$ Gaussians with the lowest importance score defined as the sum of the norm of the gradient of all parameters of each Gaussian:
@@
where, with some abuse of notation, $\texttt{lowest}_{k_{l}}$ denotes the set of the $k_{l}$ Gaussians with the lowest importance. Note that the difference $|\mathcal{G}_{l}|-|\mathcal{G}_{l-1}|$ is our $k_{l}$. The selected Gaussians will be associated with $\mathcal{E}_{l}$ and excluded (masked) from the next iterations. We reset the gradients and repeat the process. At the end of this iterative process, important Gaussians (with high gradient values) are placed at lower layers, while less important Gaussians are assigned to higher levels, where they model finer details and geometries. Although this phase requires $L-1$ steps, it tends to be very fast (1-2 minutes in total to build the hierarchy) as it only requires a single pass through the training set per step.
3.2 Fine-Tuning and SHs culling
This phase involves $F$ fine-tuning iterations designed to improve reconstruction quality and enforce a smooth transition between levels of detail. A random level $l$ is uniformly sampled at each iteration and a fine-tuning iteration is performed only on the $\mathcal{G}_{l}$ Gaussians associated with the selected level $l$. This optimization further improves reconstruction quality, especially at the lower levels. During fine-tuning, Gaussian parameters are reduced by selectively culling the spherical harmonics. Following the approach proposed by [31], SH coefficients are analyzed halfway through the fine-tuning phase to identify which primitives can be represented with 0, 1, 2, or 3 bands. The SH function is evaluated across all input views to determine whether a primitive requires view-dependent effects, based on the extent of color variation between views. If a primitive maintains a consistent color across all views, view-dependent effects are considered unnecessary. In addition, higher-order bands can be discarded if the color remains stable with lower-order SH coefficients. This phase significantly reduces the model’s memory cost (with negligible quality loss) since each Gaussian is represented by 59 32-bit floating point numbers, of which 45 are dedicated to the SH coefficients.
3.3 Scalable Compression
After fine-tuning, the compression step is performed. Compression is applied independently to the first LoD $\mathcal{G}_{0}$ and the following enhancements $\mathcal{E}_{l}$. Initially, a codebook is generated for each parameter (excluding the position of the primitives) using $K$-means clustering. In particular, following [31], codebooks of 256 entries are created for the opacity, the three scaling components, the real and imaginary parts of the rotation quaternion, the base color coefficients, and one for the 3-channel color component for each of the 15 SH coefficient groups. Codebook entries and positions are further quantized to 16-bit half-precision floats: in this way, we have achieved scalability since every LoD can be progressively decoded starting from the base one. It is worth pointing out that we move along the rate/distortion curve by just concatenating further enhancements. As shown in Sec. 4, our results are very competitive with the recent state-of-the-art, with the added value of a single scalable model obtained with a single training process. In conclusion, GoDe achieves both scalable compression and progressive LoD. Scalable compression can be exploited in many application scenarios beyond the simple memory-saving result, e.g. adaptive HTTP streaming [32]. The hierarchical organization of GoDe enhancements can be exploited for adaptive rendering at different LoD: specifically, the enhancements can be progressively splatted until a given target criterion, e.g. minimum frame rendering rate, is met. If rendering speed is the only goal, the GoDe compression phase can be skipped.

### v2

@@ Abstract
3D Gaussian Splatting enhances real-time performance in novel view synthesis by representing scenes with mixtures of Gaussians and utilizing differentiable rasterization. However, the high storage and VRAM requirements of 3DGS limit its deployment in resource-constrained environments, necessitating effective pruning and compression strategies to enable broader applicability. While existing methods are effective in certain scenarios, they lack scalability and adaptability to varying computational constraints, such as hardware capabilities or bandwidth limitations, often necessitating costly model retraining under different configurations. In this work, we introduce a novel and model-agnostic scalable compression framework for 3D Gaussian Splatting models. By leveraging our gradient-informed masking and compression algorithm, we construct a multi-level Gaussian hierarchy that efficiently balances quality and storage efficiency. At inference time, our method enables dynamic scaling across multiple compression levels without retraining, achieving comparable or superior results to state-of-the-art non-scalable compression approaches. We validate our approach on typical datasets and benchmarks, showcasing low distortion and substantial gains in terms of scalability and adaptability. The source code, provided in the supplementary material, will be open-sourced upon acceptance of the paper.
@@ 1 Introduction
Figure 1: Our method builds a progressive Gaussian hierarchy, enabling adaptive LoD. Combined with compression, it supports multiple rates (eight in the figure), from high to low compression. The model can dynamically adjust based on criteria such as a target minimum FPS.
@@
To address this scalability issue, we propose GoDe (Gaussians on Demand), a novel approach that enables scalable compression. As shown in Fig. 1, GoDe enables a 3DGS model to scale across different compression rates, without the need for retraining specialized models. Specifically, GoDe employs a progressive Level of Detail (LoD) approach: Gaussians are organized into discrete and progressive levels, with each level adding more detail (enhancement) to the previous one. This progressive structure allows 3DGS models to be highly scalable across different scenarios. In fact, by adding or removing enhancement layers, we can easily adjust the trade-off between quality and performance. When combined with a compression technique scheme, this approach enables scalable compression. Each level of detail can be seen as a distinct compression level, where higher LoDs corresponds to lower compression rate and lower LoDs corresponds to a higher compression rate. This allows for navigating the rate/distortion curve by simply adding or removing enhancement layers. To build the hierarchy, inspired by iterative gradient-informed pruning approaches, we developed an iterative gradient-informed masking methods that progressively constructs the hierarchy in a top-down manner (from high to low levels). This methods is general and compatible with popular 3DGS baselines. A quantization-aware fine-tuning phase ensures coherence and smoothness across the levels, while a final compression step further reduce model size. In summary, our main contributions are:
@@
  • A general method for progressive LoD (Sec. 3.1): We propose a novel gradient-based iterative masking technique, which effectively organizes primitives into a hierarchy of progressive levels. Our approach demonstrates high scalability (Sec. 4.4.1) and outperforms other methods in organizing the hierarchy (Sec. 4.4.2).
  • Smooth and continuous transitions between levels (Sec. 3.2): Our fine-tuning procedure provides an effective method for achieving smoother transitions, significantly enhancing reconstruction quality by finding the optimal balance between different levels of detail. Transitions are further smoothed through a linear opacity interpolation mechanism, minimizing pop-up effects (Sec. 4.3).
  • Scalable Compression (Sec. 3.3): By incorporating quantization-aware fine-tuning and entropy coding, our method achieves state-of-the-art compression-quality trade-offs (Sec. 4.2), enabling navigation along the rate-distortion curve. This results in a negligible quality loss compared to non-compressed models (Sec. 4.4.3).
@@
We evaluate our method on various 3DGS architectures and compare it against the most recent state-of-the-art methods, showing comparable and even superior performances than non-scalable method. Our method allows models to scale across a wide range of compression levels, from low to very high, achieving up to 99.37% reduction in model size and 98.36% in the number of primitives in the highest compression rate with respect to non-compressed models, while preserving comparable full-size model quality at the lowest compression rate.
@@ 2 Background and Related Work
LoD is a well-established technique in computer graphics, used to optimize performance and improve visual quality. Recent work has extended LoD to 3DGS, primarily to mitigate aliasing artifacts [41] and improve performance, especially for large-scale datasets [24]. These methods typically involve learning multiple levels of detail within the scene representation, followed by selective rendering according to heuristics such as pixel coverage (i.e., the number of pixels occupied by a Gaussian projection on the image plane) and distance from the camera. Other works propose a hierarchical representation, often structured as a tree, as proposed by [34, 19]. This method dynamically implements LoD by selectively rendering Gaussians up to a specific level within the structure based on certain heuristics. Our technique is orthogonal to these works and can be integrated into dynamic LoD systems, where levels serve as an upper bound on active primitives. A related resolution-based LoD method [35] iteratively refines from a minimal representation, but suffers from three key limitations: high iteration count (30k per level), dependence on stored opacity values from previous levels, and rapid Gaussian proliferation at higher levels. In contrast, our approach supports an arbitrary number of levels without additional overhead, requiring only 30k iterations in total while maintaining a compact parameter count. Moreover, it generalizes across models and integrates compression for further efficiency.
@@
Since the introduction of hybrid radiance fields models that incorporate auxiliary data structures to accelerate training and inference times, such as voxel grids [36, 6, 14, 30], research has focused on compression [9] and complexity reduction [10]. For instance, [9] demonstrates how an iterative-magnitude pruning approach is highly effective in reducing the number of parameters even in the context of radiance fields. These techniques have recently been adapted in the field of 3DGS. In fact, despite its advantages, 3DGS is limited by high storage and memory demands, reducing its application, for example, on low-end hardware. Current compression methods aim at achieving the highest quality at minimal storage size. Typically, compact representations involve post hoc Gaussian pruning followed by compression stages, including quantization (both during and after training) and entropy coding [2, 1, 32, 39]. Some methods operate directly during training, with the aim of limiting the number of Gaussians through learned masks and optimized encoding schemes [15, 7, 23, 33, 37]. Other studies enhance the representation by employing neural networks for on-the-fly Gaussian generation [25], tri-planes [38], or by using auxiliary structures such as hash grids to store spherical harmonics [7, 22], further reducing storage demands. In particular, Scaffold-GS [25] reduces storage by up to 5 times by replacing Gaussians with anchor points. These anchor points, which are significantly fewer in number, are associated with a set of features, offsets, and scalings. A set of shallow neural networks processes each anchor point and generates the parameters of view-adaptive Gaussians. However, while these methods enable compression, they lack practical scalability. Adjusting compression rates requires either retraining from scratch or additional fine-tuning for each level, limiting their adaptability to scenarios with varying hardware constraints or bandwidth limitations. In contrast, our method incorporates a scalable compression scheme, enabling multiple compression levels from a single training run. This allows instant navigation of the rate-distortion curve without additional optimization. Moreover, our approach achieves performance comparable to, or even exceeding, dedicated non-scalable compression techniques.
@@ 3 Gaussians on demand
Figure 2: Graphical overview of our method GoDe. 1) Starting from a pre-trained model, we build a progressive hierarchy of Gaussians through iterative gradient-informed masking. 2) A quantization-aware fine tuning stage enhances the model’s quality and reduces its memory footprint. 3) The model is further compressed through entropy coding.
This section presents our framework for progressive LoD and scalable compression. Our approach extends the standard 3DGS framework by introducing a progressive hierarchy of Gaussians, with varying levels of detail. The objective is to develop a single, scalable model that can dynamically adapt to different compression levels, rendering budgets, and hardware or network constraints. We present a novel algorithm that constructs this hierarchy implicitly, enabling the model to learn multiple levels of detail. We first describe how to build the progressive hierarchy of Gaussians from a pre-trained 3DGS model (Sec.3.1), which forms the basis for constructing adaptable LoDs. Then, we explain how to train and fine-tune Gaussians for each level of detail (Sec. 3.2). Finally, we describe our scalable compression method (Sec.  3.3).
@@ 3.1 Hierarchy of Gaussians
$\mathcal{G}=(\mathcal{G}_{0},\leavevmode\nobreak\ \mathcal{E}_{1},\dots,\mathcal{E}_{l},\dots,\mathcal{E}_{L-1}),$ (1)
@@
$\mathcal{G}_{0}\cap\mathcal{E}_{1},\dots,\cap\leavevmode\nobreak\ \mathcal{E}_{l},\dots,\cap\leavevmode\nobreak\ \mathcal{E}_{L-1}=\emptyset.$ (3)
@@
and $c_{\text{min}}=|\mathcal{G}_{0}|$ is the number of Gaussians at the first LoD level (chosen as desired). We choose this progression since the rate-distortion functions generally follow a logarithmic law (see the theoretical rate-distortion curve for a Gaussian source [8]). This logarithmic progression is chosen to reflect the diminishing returns of adding additional Gaussians, but alternative progression schemes could be employed based on specific application constraints (e.g., computational capabilities of the rendering devices). To organize the Gaussians of the model into a hierarchical structure, we employ an iterative gradient-informed masking process. This approach follows a top-down strategy, starting with the construction of the higher enhancement levels first, namely, progressively building the model’s hierarchy from layer $l=L-1$ to layer $l=1$. The base layer $\mathcal{G}_{0}$ will consist of the remaining Gaussians at the end of the process. The gradient of each Gaussian parameter serves as a critical indicator of its relevance to the overall model. A Gaussian with a low gradient indicates minimal contribution to the loss function. Conversely, a high gradient suggests that the Gaussian is essential and has a significant impact on the loss. We leverage this gradient information to construct the hierarchy. Specifically, to create the enhancement set $\mathcal{E}_{l}$, we accumulate gradients on $\mathcal{G}_{l}$ using the training set in order to select the $k_{l}$ Gaussians with the lowest importance score defined as the sum of the norm of the gradient of all parameters of each Gaussian:
@@
where $\texttt{lowest}_{k_{l}}$ selects the set of the $k_{l}$ Gaussians with the lowest gradient norm. Note that the difference $|\mathcal{G}_{l}|-|\mathcal{G}_{l-1}|$ is our $k_{l}$. The selected Gaussians are associated with $\mathcal{E}_{l}$ and masked from further consideration in the next iterations. We reset the gradients and repeat the process. Although this phase requires $L-1$ steps, it tends to be very fast (1-2 minutes in total to build the hierarchy) as it only requires a single pass through the training set per step. By the end of this process, important Gaussians (those with high gradients) are positioned at lower levels, contributing to coarse details, while less important Gaussians are placed at higher levels, modeling finer geometries and details.
3.2 Quantization-Aware Fine-tuning
This phase involves $F$ fine-tuning iterations. A random level $l$ is uniformly sampled at each iteration and a fine-tuning iteration is performed only on the $\mathcal{G}_{l}$ Gaussians associated with the selected level $l$. This optimization enhances reconstruction quality, particularly at the lower levels, while also ensuring smoother transitions between them. During fine-tuning, we integrate quantization-aware training by quantizing spherical harmonic coefficients, opacity, scaling, and rotation parameters. Quantization is applied before the forward pass. To enable gradient flow through quantized parameters, a straight-through estimator is used, which allows backpropagation to proceed as if quantization were absent. More details are given in the Supplementary Sec. 7. Since spherical harmonic coefficients are the most storage-intensive parameters, we further apply a L1 loss to encourage sparsity. Specifically, given the set of spherical harmonics coefficients $\mathcal{C}_{l}$ associated with the Gaussians in level $\mathcal{G}_{l}$ we introduce the penalty term:
$\mathcal{L}_{\text{SH}}=\lambda\sum_{i}\|\mathbf{C}_{l,i}\|_{1}.$ (7)
3.3 Scalable Compression and Adaptive Rendering
After fine-tuning, a further compression step is applied. Compression is performed on the first LoD $\mathcal{G}_{0}$ as well as on the subsequent enhancement levels $\mathcal{E}_{l}$. This inherently enables scalable compression, where each level of detail can be interpreted as a compression stage. Specifically, lower LoDs correspond to higher compression rates, whereas higher LoDs yield lower compression. Each level can be progressively decoded starting from the previous one. Thus, the rate-distortion trade-off can be controlled simply by concatenating enhancement layers. As demonstrated in Sec. 4, our approach achieves highly competitive results compared to the state-of-the-art, with the added benefit of relying on a single scalable model derived from a unified training process. This hierarchical organization can also be leveraged for adaptive rendering at varying levels of detail. More specifically, enhancement layers can be progressively splatted until a target criterion, such as a minimum frame rendering time threshold, is met. To prevent pop-up artifacts and ensure smooth transitions between LoDs, we introduce continuous transitions by linearly interpolating the opacity values of Gaussians in the subsequent enhancement layer. We define a vector of $N$ increasing interpolation factors:
$\mathbf{t}=\{t_{1},t_{2},\dots,t_{N}\},\quad t_{i}\in[0,1],\quad t_{1}<t_{2}<\dots<t_{N}$ (8)
and scale the Gaussians’ opacities in the enhancement layer accordingly. In conclusion, GoDe enables 3D Gaussian Splatting models to achieve scalable compression and adaptive rendering, providing state-of-the-art rate-distortion trade-offs.